# ClusterRCA: Network Failure Diagnosis in HPC Systems Using Multimodal Data

### v1

@@ IV-E Evaluation on Scalability (RQ4)
In real HPC systems, the dynamic allocation of computing resources by schedules determines the number and specific nodes for each job of HPC applications. Therefore, the failure diagnosis method must be feasible in practice and consider the HPC cluster’s dynamic topology, ensuring that the trained model exhibits robust scalability. Specifically, the failure diagnosis model trained on a small-scale cluster can also be directly used for a larger-scale cluster failure diagnosis while ensuring a certain accuracy. Due to the limitation of the experimental environment, we simulate two datasets of 16 nodes and 32 nodes, respectively, namely 16RANK and 32RANK (A is used for training, B is used for testing). As listed in Table IV, we test ClusterRCA on 16RANK_B and 32RANK_B, respectively, and use the small-scale data (16RANK_A) to train ClusterRCA and evaluate its failure diagnosis ability on the larger-scale data (32RANK_B). The results show that ClusterRCA has $Avg@5$ greater than 0.93 on both scales. The $Avg@5$ of ClusterRCA trained on 16RANK_A is very close to that of ClusterRCA trained on 32RANK_A (0.93 & 0.94). The experiment proves that ClusterRCA also has excellent scalability.
@@
IV-F Threat to Validity
Dataset limitations. Compared to complex HPC systems, our cluster size is relatively small. Due to signing a confidentiality agreement with $\boldsymbol{A}$ company, we cannot open-source the original data of the experiment. Despite efforts to localize publicly available HPC-related network failure datasets, the scarcity of research in HPC failure diagnosis thwarted our endeavors. However, based on our experiment, ClusterRCA proves effective and robust. Its applicability extends to larger HPC systems and more intricate failure scenarios, presenting promising prospects.
Data modality limitations. Our work involves using three modalities of data (i.e., metrics, logs, and topology), while some HPC systems may lack metric or log collection. Since the feature extraction of metrics and logs is loosely coupled, ClusterRCA can still function normally in the absence of a particular data modality.
@@ V Discussion
V-A Lessons Learned
Clock synchronization. When a failure occurs, it propagates from the culprit node to other nodes, causing different patterns of failure to appear on different types of nodes. If the clock difference between different nodes is too large, it may cause the key failure feature of this job to be lost. For example, the clock of node A is relatively delayed by $\delta$, and the cluster starts to fail at $t_{a}$ (the time recorded by A at $t_{a}-\delta$), while the time range for the diagnostic data is from $t_{b}$ to $t_{c}$. If $t_{a}-\delta$ is earlier than $t_{b}$, then the failure feature of node A will be missed, and thus affect the subsequent construction of the failure propagation graph and inference. Therefore, the computing nodes and switch nodes in the cluster need to use clock synchronization algorithms to ensure that the monitoring data are aligned as much as possible.
Failure injection. In our experiment, we inject failure by running scripts, and then construct corresponding failure labels (culprit node and failure type) for final performance evaluation. Then, due to some unknown reasons, sometimes failure injection may fail, without causing HPC performance degradation. In this case, the failure labels are incorrect. To ensure the accuracy of evaluation results, we will filter out the samples of incorrect failure injection according to certain rules and human verification.
@@ V-B Threat to Validity
Data set limitations. Compared to complex HPC systems, our cluster size is relatively small. Due to signing a confidentiality agreement with $\boldsymbol{A}$ company, we cannot open-source the original data of the experiment. Although we also tried to find public network failure data sets related to HPC, we failed in the end due to the scarcity of HPC failure diagnosis research work. However, according to our experiment, ClusterRCA is effective and robust. It can also be applied effectively to larger HPC systems and more complex failure situations, which is very promising.
Data modalities limitations. Our work involves the use of three multimodal monitoring data (metric, log, topology), while some HPC systems may lack metric or log collection. Since the feature extraction of metrics and logs are in a loose coupling way, ClusterRCA can still work normally in a data modality missing.
@@ VI Related Work
Graph-based diagnostic methods. Graph-based diagnostic methods establish directed relationships between cluster nodes based on topology or system flow data. They then use random walks or causal inference to find culprit nodes. For example, MicroCause[25] builds a directed acyclic graph showing metric dependencies using the PC algorithm and localizes the culprit node with an improved random walk. Similar methods is AutoMap[19]. However, the PC algorithm must be improved to find cross-node metric propagation with multiple similar nodes. Sieve[23] uses clustering to extract representative metrics for each node, combines business relationships into node dependency graphs, and makes causal inferences to determine the culprit node. However, HPC nodes lack business relationships. CauseInfer[24] automatically builds node dependency graphs from network traffic data between monitored nodes and uses the PC algorithm for metric causal graphs within nodes. It then makes a two-layer causal inference to identify the culprit metric. Nonetheless, due to the high-performance computing demands, monitoring HPC network traffic at a fine granularity is not feasible.
Combined Classifier and Graph diagnostic methods. CloudRCA[3] from Alibaba combines metrics and logs. It uses the PC algorithm to associate cluster features with failure types and a hierarchical Bayesian model to classify failure types but not specific culprit nodes. ART [39] designs a unified fault representation and automatically identifies anomalies based on extreme value theory thresholds, which is combined with cut-tree clustering to achieve unsupervised fault type triage, and cosine similarity to accomplish root cause localization. Eadro [40] jointly trains anomaly detection and root cause localization models through shared representations, and the loss function optimizes both tasks simultaneously, effectively using shared knowledge to reduce the negative impact of detection errors on localization.
@@ VII Conclusion
Network failure diagnosis is of great importance for sustaining HPC system operability. In this paper, based on previous cluster failure diagnosis research, we first elucidate the challenges of culprit node localization and failure type identification from the heterogeneous data in HPC systems. We then propose a network failure diagnosis system, ClusterRCA. It leverages NIC-centered feature engineering to integrate metrics and logs from topologically adjacent HPC components, addressing the challenge of data diversity and device heterogeneity. Its failure diagnosis algorithm amalgamates classifier-based and graph-based methods so ClusterRCA can construct a directed graph to localize the culprit nodes accurately and establish a highly accurate classifier to determine failure types. Finally, we evaluate ClusterRCA on experimental datasets collected by a top-tier global HPC device vendor ($\boldsymbol{A}$ company). All results demonstrate its robust diagnostic capabilities.

### v2

@@ IV-E Evaluation on Scalability (RQ4)
In real HPC systems, the dynamic allocation of computing resources by schedules determines the number and specific nodes for each job of HPC applications. Therefore, the failure diagnosis method must be feasible in practice, and the HPC cluster’s dynamic topology must be considered, ensuring that the trained model exhibits robust scalability. Specifically, the failure diagnosis model trained on a small-scale cluster can also be directly used for a larger-scale cluster failure diagnosis while ensuring a certain accuracy. We collected two datasets with 16 and 32 nodes, respectively, in an HPC system simulation environment, and named them 16RANK and 32RANK (A is used for training, B is used for testing). As listed in Table IV, we test ClusterRCA on 16RANK_B and 32RANK_B, respectively, and use the small-scale data (16RANK_A) to train ClusterRCA and evaluate its failure diagnosis ability on the larger-scale data (32RANK_B). The results show that ClusterRCA has $Avg@5$ greater than 0.93 on both scales. The $Avg@5$ of ClusterRCA trained on 16RANK_A is very close to that of ClusterRCA trained on 32RANK_A (0.93 & 0.94). The experiment proves that ClusterRCA also has excellent scalability.
@@ V Discussion
V-A Observed Issues and Solutions
Clock synchronization. When a failure occurs, it propagates from the culprit node to other nodes, causing different failure patterns to appear on different types of nodes. If the clock difference between different nodes is too large, it may cause the key failure feature of this job to be lost. For example, the clock of node A is relatively delayed by $\delta$, and the cluster starts to fail at $t_{a}$ (the time recorded by A at $t_{a}-\delta$), while the time range for the diagnostic data is from $t_{b}$ to $t_{c}$. If $t_{a}-\delta$ is earlier than $t_{b}$, then the failure feature of node A will be missed, and thus affect the subsequent construction of the failure propagation graph and inference. Therefore, the computing nodes and switch nodes in the cluster need to use clock synchronization algorithms to ensure that the monitoring data is aligned as much as possible.
Failure injection. In our experiment, we inject failure by running scripts, and then construct corresponding failure labels (culprit node and failure type) for final performance evaluation. Then, for some unknown reason, sometimes failure injection may fail, without causing HPC performance degradation. In this case, the failure labels are incorrect. To ensure the accuracy of evaluation results, we will filter out the samples of incorrect failure injection according to specific rules and human verification.
@@ V-B Threat to Validity
Data set limitations. Compared to complex HPC systems, our cluster size is relatively small. Due to signing a confidentiality agreement with the $\boldsymbol{A}$ company, we cannot open-source all the original data of the experiment, as well as the fault injection scripts that involve system structure information. Although we also tried to find public network failure data sets related to HPC, we failed in the end due to the scarcity of HPC failure diagnosis research work. However, according to our experiment, ClusterRCA is effective and robust. It can also be applied effectively to larger HPC systems and more complex failure situations, which is very promising.
Data modalities limitations. Our work involves using three multimodal monitoring data (metric, log, topology), while some HPC systems may lack metric or log collection. Since the feature extraction of metrics and logs is loosely coupled, ClusterRCA can still work normally in a data modality that is missing.
@@ VI Related Work
Graph-based diagnostic methods. Graph-based diagnostic methods establish directed relationships between cluster nodes based on topology or system flow data. They then use random walks or causal inference to find culprit nodes. For example, MicroCause[13] builds a directed acyclic graph showing metric dependencies using the PC algorithm and localizes the culprit node with an improved random walk. A similar method is AutoMap[12]. However, the PC algorithm must be improved to find cross-node metric propagation with multiple similar nodes. Sieve[10] uses clustering to extract representative metrics for each node, combines business relationships into node dependency graphs, and makes causal inferences to determine the culprit node. However, HPC nodes lack business relationships. CauseInfer[11] automatically builds node dependency graphs from network traffic data between monitored nodes and uses the PC algorithm for metric causal graphs within nodes. It then makes a two-layer causal inference to identify the culprit metric. Nonetheless, due to the high-performance computing demands, monitoring HPC network traffic at a fine granularity is not feasible.
Combined Classifier and Graph diagnostic methods. CloudRCA[14] from Alibaba combines metrics and logs. It uses the PC algorithm to associate cluster features with failure types and a hierarchical Bayesian model to classify failure types, but not specific culprit nodes. ART [15] designs a unified fault representation and automatically identifies anomalies based on extreme value theory thresholds, which is combined with cut-tree clustering to achieve unsupervised fault type triage, and cosine similarity to accomplish root cause localization. Eadro [16] jointly trains anomaly detection and root cause localization models through shared representations, and the loss function optimizes both tasks simultaneously, effectively using shared knowledge to reduce the negative impact of detection errors on localization.
@@ VII Conclusion
Network failure diagnosis is of great importance for sustaining HPC system operability. In this paper, based on previous cluster failure diagnosis research, we first elucidate the challenges of culprit node localization and failure type identification from the heterogeneous data in HPC systems. We then propose a network failure diagnosis system, ClusterRCA. It leverages NIC-centered feature engineering to integrate metrics and logs from topologically adjacent HPC components, addressing the challenge of data diversity and device heterogeneity. Its failure diagnosis algorithm amalgamates classifier-based and graph-based methods so ClusterRCA can construct a directed graph to localize the culprit nodes and establish a highly accurate classifier to determine failure types. Finally, we evaluate ClusterRCA on experimental datasets from a top-tier global HPC device vendor ($\boldsymbol{A}$ company). All results demonstrate its robust diagnostic capabilities.
Acknowledgment
This work is supported by the Advanced Research Project of China (No. 31511010501), the National Natural Science Foundation of China (62272249, 62302244), and the Fundamental Research Funds for the Central Universities (XXX-63253249).